OpenStudy (anonymous):

Solve 5x^2 + 3x - 4 = 0 by completing the square.

OpenStudy (owlcoffee):

So we know that: \[(a+b)^{2}=a ^{2}+2ab+b ^{2}\] Now, looking at the problem,. We need to find a number that can help us complete the formula.

OpenStudy (dumbcow):

\[5(x^{2}+\frac{3}{5}x) = 4\] \[x^{2}+\frac{3}{5}x = \frac{4}{5}\] \[x^{2}+\frac{3}{5}x +(\frac{3}{10})^{2} = \frac{4}{5}+(\frac{3}{10})^{2}\] \[(x+\frac{3}{10})^{2} = \frac{4}{5}+\frac{9}{100}\] \[(x+\frac{3}{10})^{2} = \frac{89}{100}\] \[x + \frac{3}{10} = \frac{\pm \sqrt{89}}{10}\] \[x = \frac{-3 \pm \sqrt{89}}{10}\]

OpenStudy (anonymous):

thank you guys